OpenStudy (anonymous):

If n is a term of the sequence 2, 5, 8, 11, ..., which expression would give the value of n? 2n+ 1 2n- 1 3n- 1 n+ 3

OpenStudy (anonymous):

Think about this. What adds to 2 to get 5, as well as getting from 5 to 8?

OpenStudy (anonymous):

3n-1

OpenStudy (anonymous):

3

OpenStudy (anonymous):

3n-1

OpenStudy (anonymous):

Yep, 3. So, n + 3 would be the correct answer here.

OpenStudy (anonymous):

3n-1 e.g. 3(2)-1=5 this means there is a difference of 3 in each number

OpenStudy (anonymous):

11 = 8(3) - 1 11 = 24 - 1 11 is not 23, so it doesn't work.